\begin{thebibliography}{10}

\bibitem{ALLAIRE-02}
Gr\'{e}goire Allaire.
\newblock Shape optimization by the homogenization method.
\newblock 146:xvi+456, 2002.

\bibitem{BBM-01}
Jean Bourgain, Haim Brezis, and Petru Mironescu.
\newblock Another look at {S}obolev spaces.
\newblock pages 439--455, 2001.

\bibitem{BRASCO-PARINI-16B}
Lorenzo Brasco and Enea Parini.
\newblock The second eigenvalue of the fractional {$p$}-{L}aplacian.
\newblock {\em Adv. Calc. Var.}, 9(4):323--355, 2016.

\bibitem{BRASCO-PARINI-16}
Lorenzo Brasco, Enea Parini, and Marco Squassina.
\newblock Stability of variational eigenvalues for the fractional
  {$p$}-{L}aplacian.
\newblock {\em Discrete Contin. Dyn. Syst.}, 36(4):1813--1845, 2016.

\bibitem{BREZIS-LIBRO}
Haim Brezis.
\newblock {\em Functional analysis, Sobolev spaces and partial differential
  equations}.
\newblock Universitext. Springer, New York, 2011.

\bibitem{BUTTAZZO-BUCUR}
Dorin Bucur and Giuseppe Buttazzo.
\newblock {\em Variational methods in shape optimization problems}, volume~65
  of {\em Progress in Nonlinear Differential Equations and their Applications}.
\newblock Birkh\"{a}user Boston, Inc., Boston, MA, 2005.

\bibitem{BUTTAZZO-DALMASO-93}
Giuseppe Buttazzo and Gianni Dal~Maso.
\newblock An existence result for a class of shape optimization problems.
\newblock {\em Arch. Rational Mech. Anal.}, 122(2):183--195, 1993.

\bibitem{CERESABONDER24}
Ignacio Ceresa~Dussel and Juli\'{a}n Fern\'{a}ndez~Bonder.
\newblock Existence of eigenvalues for anisotropic and fractional anisotropic
  problems via ljusternik-schnirelman theory.

\bibitem{CERESABONDER23}
Ignacio Ceresa~Dussel and Juli\'{a}n Fern\'{a}ndez~Bonder.
\newblock A {B}ourgain-{B}rezis-{M}ironescu formula for anisotropic fractional
  {S}obolev spaces and applications to anisotropic fractional differential
  equations.
\newblock {\em J. Math. Anal. Appl.}, 519(2):Paper No. 126805, 25, 2023.

\bibitem{CHAKER-23}
Jamil Chaker, Minhyun Kim, and Marvin Weidner.
\newblock The concentration-compactness principle for the nonlocal anisotropic
  {$p$}-{L}aplacian of mixed order.
\newblock {\em Nonlinear Anal.}, 232:Paper No. 113254, 18, 2023.

\bibitem{DALMASO-93}
Gianni Dal~Maso.
\newblock {\em An introduction to {$\Gamma$}-convergence}, volume~8 of {\em
  Progress in Nonlinear Differential Equations and their Applications}.
\newblock Birkh\"{a}user Boston, Inc., Boston, MA, 1993.

\bibitem{BONDER-RITTORTO-18}
Juli\'{a}n Fern\'{a}ndez~Bonder, Antonella Ritorto, and Ariel~Martin Salort.
\newblock A class of shape optimization problems for some nonlocal operators.
\newblock {\em Adv. Calc. Var.}, 11(4):373--386, 2018.

\bibitem{BONDER-SALORT-20}
Julian Fern\'{a}ndez~Bonder and Ariel Salort.
\newblock Stability of solutions for nonlocal problems.
\newblock {\em Nonlinear Anal.}, 200:112080, 13, 2020.

\bibitem{FRANZINA-14}
Giovanni Franzina and Giampiero Palatucci.
\newblock Fractional {$p$}-eigenvalues.
\newblock {\em Riv. Math. Univ. Parma (N.S.)}, 5(2):373--386, 2014.

\bibitem{FUSCO-19}
Nicola Fusco, Shirsho Mukherjee, and Yi~Ru-Ya Zhang.
\newblock A variational characterisation of the second eigenvalue of the
  {$p$}-{L}aplacian on quasi open sets.
\newblock {\em Proc. Lond. Math. Soc. (3)}, 119(3):579--612, 2019.

\bibitem{LINDQVIST-14}
Erik Lindgren and Peter Lindqvist.
\newblock Fractional eigenvalues.
\newblock {\em Calc. Var. Partial Differential Equations}, 49(1-2):795--826,
  2014.

\bibitem{MOTREANU14}
Dumitru Motreanu, Viorica~Venera Motreanu, and Nikolaos Papageorgiou.
\newblock {\em Topological and variational methods with applications to
  nonlinear boundary value problems}.
\newblock Springer, New York, 2014.

\bibitem{PONCE-04}
Augusto~C. Ponce.
\newblock A new approach to {S}obolev spaces and connections to
  {$\Gamma$}-convergence.
\newblock {\em Calc. Var. Partial Differential Equations}, 19(3):229--255,
  2004.

\bibitem{SOKOLOWSKI-ZOLESIO}
Jan Soko\l~owski and Jean-Paul Zol\'{e}sio.
\newblock {\em Introduction to shape optimization}, volume~16 of {\em Springer
  Series in Computational Mathematics}.
\newblock Springer-Verlag, Berlin, 1992.
\newblock Shape sensitivity analysis.

\bibitem{WARMA-15}
Mahamadi Warma.
\newblock The fractional relative capacity and the fractional {L}aplacian with
  {N}eumann and {R}obin boundary conditions on open sets.
\newblock {\em Potential Anal.}, 42(2):499--547, 2015.

\end{thebibliography}

\documentclass{amsart}

\usepackage{amsfonts}
\usepackage{color}
\usepackage{graphicx}
\usepackage[bookmarksnumbered,colorlinks, linkcolor=blue, citecolor=red, pagebackref, bookmarks, breaklinks]{hyperref}
\usepackage{enumitem}
\usepackage[bookmarksnumbered,colorlinks, linkcolor=blue, citecolor=red, pagebackref, bookmarks, breaklinks]{hyperref}
\usepackage{hyperref}

\newtheorem{thm}{Theorem}[section]

\newtheorem{lema}[thm]{Lemma}
\newtheorem{prop}[thm]{Proposition}

\theoremstyle{definition}
\newtheorem{defn}[thm]{Definition}
\theoremstyle{remark}
\newtheorem{rem}[thm]{Remark}

\numberwithin{equation}{section}
\newcommand{\A}{\mathcal A}
\newcommand{\R}{\mathbb R}
\newcommand{\N}{\mathbb N}

\newcommand{\s}{\mathbf{s}}
\newcommand{\p}{\mathbf{p}}
\newcommand{\K}{\mathcal K}
\newcommand{\D}{\mathcal{D}}

\def\supp{\mathop{\text{\normalfont supp}}}
\def\diver{\mathop{\text{\normalfont div}}}

\begin{document}
	\title[Shape optimization problems]{Shape optimization problems involving nonlocal and nonlinear operators.}
	
	\author[Ignacio Ceresa Dussel]{Ignacio Ceresa Dussel}
	
	\address{Instituto de C\'alculo, CONICET\\
		Departamento de Matem\'atica, FCEN - Universidad de Buenos Aires\\
		Ciudad Universitaria, 0+$\infty$ building, C1428EGA, Av. Cantilo s/n\\
		Buenos Aires, Argentina}
	
	\begin{abstract}
	In this research, we investigate a general shape optimization problem in which the state equation is expressed using a nonlocal and nonlinear operator.
	
	We prove the existence of a minimum point for a functional $F$ defined on the family of all 'quasi-open' subsets of a bounded open set $\Omega$ in $\mathbb{R}^n$. This is ensured under the condition that $F$ demonstrates decreasing behavior concerning set inclusion and is lower semicontinuous with respect to a suitable topology associated with the fractional $p$-Laplacian under Dirichlet boundary conditions.
	
	Moreover, we study the asymptotic behavior of the solutions when $s\to1$ and extend this result to the anisotropic case.
	
	\end{abstract}
	
	\subjclass[2020]{ 35R11, 49Q99 }
	
	
	\keywords{Shape optimization, fractional Laplacian, $\Gamma$-convergence, Anisotropic and Sobolev spaces}

	\maketitle

\section{Introduction}
In this study, we will address shape optimization problems. These type of problems have been extensively considered in the literature. Only to mention some references, we refer the reader to the books Bucur and Buttazzo \cite{BUTTAZZO-BUCUR}, Allaire \cite{ALLAIRE-02} and Sokolowski and Zolésio \cite{SOKOLOWSKI-ZOLESIO}.

 In their most general form, these problems involve a \emph{cost functional} $F$ and a class of \emph{admissible domains} $\A$. The objective is to solve the minimization problem stated as
\begin{equation}\label{problem}
	\min_{A\in\A}F(A).
\end{equation}
Issues of this nature naturally emerge in structural optimization, particularly when one wants to determine the optimal design of a system governed by partial differential equations.

In many of the existing literature sources, the cost functional $F$ is typically formulated in terms of a function $u_A$, which serves as the solution of a state equation to be computed over $A$ in integral form. Additionally, these problems often include constraints on the Lipschitz constant of $\partial A$. However, such constraints have sometimes been viewed as unnatural in practical applications. Consequently, in their research \cite{BUTTAZZO-DALMASO-93}, the authors introduced the family $\A$ of admissible domains as all \emph{quasi open} subsets of $\Omega$ with the sole constraint of volume, $|A| = c$. With regards to the function $F$, they posit that it decreases concerning set inclusion and is lower semicontinuous with respect to a suitable topology on $\A$. To be more explicit, the relationship of $F$ with a set $A$ was as follows, \emph{ a sequence $\{A_k\}_{k\in \N}$ converges to $A$ if for every $f\in H^{-1}(\Omega)$ the solutions of the Dirichlet problems 
$$
-\Delta u_k=f \text{ in }A_k,\quad u_k=0 \in \partial A_k,
$$ converge as $k\to \infty$ to the solution of the corresponding problem in $A$.}
Therefore, the existence problem for the "linear case" was solved.

On one hand, in \cite{BONDER-RITTORTO-18}, the authors study the \emph{non-local}
linear problem. The problem was considered with the aim of minimizing, taking into account the non-local Dirichlet problems
$$
(-\Delta)^su=f \text{ in }A_k,\quad u_k=0 \in \partial A_k,
$$ 
where $-(\Delta)^s$ is the \emph{fractional Laplacian} and $f$ belong in a suitable space. 

On the other hand, in \cite{FUSCO-19}, the authors proof similar results when the partial differential equation is the local non linear $p$-Laplacian
$$
-\Delta_pu:=\diver(|\nabla u|^{p-2}\nabla u).
$$
Even more, they prove a minimax characterization of the second eigenvalue of the $p$-Laplacian operator on $p$-quasi open sets.

In this study, our initial focus will be on the natural subsequent step, which involves minimizing problems similar to \ref{problem}, where the relation between the cost functional and the domain is the non-local and non-linear Dirichlet problem
$$
(-\Delta_p)^su=f \text{ in }A_k,\quad u_k=0 \in \partial A_k,
$$ 
where the operator is the \emph{fractional $p$-laplacian}
$$
(-\Delta_p)^su:=\frac{1-s}{K(n,p)}p.v\int_{\R^n}\frac{|u(x)-u(y)|^{p-2}(u(x)-u(y))}{|x-y|^{n+sp}}\,dy.
$$ 
 Keep in mind that the fractional $p$-Laplacian is the non-local equivalent of the $p$-Laplacian. Based on \cite{BBM-01,BONDER-SALORT-20,BONDER-RITTORTO-18}, we examine the transition from non-local to local operators, specifically studying the asymptotic behavior as $s$ approaches 1.

Finally, based on \cite{CERESABONDER23,CHAKER-23}, we will introduce the anisotropic counterparts of the fractional $p$-laplacian and the $p$-laplacian, namely the \emph{pseudo $p$-laplacian}
$$
-\widetilde{\Delta}_\p u:=\sum_{i=1}^{n}(|u_{x_i}|^{p_i-2}u_{x_i})_{x_i},
$$

and the \emph{ fractional pseudo $p$-laplacian}
\begin{equation}\label{fractiona pseudo p}
(-\widetilde{\Delta}_\p)^\s u:=\sum_{i=1}^{n}(1-s_i)p.v\int_{\R}\frac{|u(x+he_i)-u(x)|^{p_i-2}(u(x+he_i)-u(x))}{|h|^{1+s_ip_i}}\,dh.
\end{equation}
In this expression, $\s=(s_1,\cdots,s_n)$ represents the differential fractional parameters and $\p=(p_1,\cdots,p_n)$ represent different integral parameters.

The paper is structured as follows: In Section \ref{sec_preliminares}, we present all preliminaries and some useful properties. In Section \ref{sec_teo1}, we will state and prove the main result, which is based on the existence of a minimizer. In Section \ref{sec_asint} discusses the asymptotic behavior of the solution as $s$ approaches 1. Finally, in Section \ref{sec_anis} prove the existence of a minimizer when the operator exhibits anisotropic characteristics.
%
 
\section{Preliminaries}\label{sec_preliminares}
In this section, we will introduce definitions and properties pertaining to the functional framework that will be employed.

Let $0 < s < 1$ and $1 < p < \infty$ be parameters, the fractional Sobolev space is defined as
\[
W^{s, p}(\R^n) = \left\{ u \in L^p(\R^n) : \frac{u(x) - u(y)}{|x - y|^{\frac{n}{p} + s}} \in L^p(\R^n\times\R^n) \right\}.
\]
This space is endowed with the norm
$$
\|u\|_{W^{s, p}(\R^n)} = \left( \|u\|_{L^p(\R^n)}^p + [u]_{s, p} \right)^{\frac{1}{p}},
$$
where
$$
[u]_{s,p} = \left( \iint_{\R^n \times \R^n} \frac{|u(x) - u(y)|^p}{|x - y|^{n+sp}} \, dxdy \right)^{\frac{1}{p}
}.
$$
With the above norm, \(W^{s, p}(\R^n)\) is a reflexive Banach space. 
	
If we consider a bounded open set of $\R^n$, namely $\Omega$, the common practice is to take into account the closure of $C^\infty_c(\R^n)$ with respect to $\|.\|_{W^{s,p}}$. Another definition is given by
	$$
	W_0^{s, p}(\Omega)=\{u\in W^{s, p}(\R^n)\, |\, u=0 \text{ in }  \R^n\setminus\Omega \}.
	$$
	It is well known that these two definitions agree when, for example, $\partial \Omega$ is Lipschitz. Therefore, since the boundary of $\Omega$ is not of our interest, we will assume that the set $\Omega$ has a Lipschitz boundary.

When $s=1$, we recover the well-known Sobolev space $W^{1,p}_0(\Omega)$, with the classical norm $\|\nabla u\|_{L^p(\Omega)}$.
\begin{defn}
	Given $0 < s \leq 1 \leq p < \infty$, the (topological) dual space of $W^{s,p}_0(\Omega)$ will be denoted by $W^{-s,p'}(\Omega)$, and $\langle\cdot,\cdot\rangle$ is the duality map between $W^{-s,p'}(\Omega)$ and $W^{s,p}_0(\Omega)$.
	
\end{defn}

\subsection{Capacity}
 In \cite{WARMA-15} the author define the $(s,p)$-Capacity of a set $A\subset\Omega$. We will recall some useful definitions and theorems about this fractional capacity, and refer the interested reader to see \cite{WARMA-15} and their references.

\begin{defn}
	Let $\Omega \subset \mathbb{R}^n$ be an open set. Given $A \subset \Omega$, for any $0 < s \leq 1<p<\infty$, we define the $(s,p)$-Capacity of $A$ as
	$$
	\text{Cap}(s,p)(A) = \inf \left\{ [u]_{s,p} : u \in W^{s,p}_0(\Omega), u \geq 0, A \subset \{u \geq 1\}^\circ \right\}.
	$$
\end{defn}
If a property $P(x)$ holds for every $x \in E$ except for those in a set $Z \subset E$ with $\text{Cap}(s,p)(Z) = 0$, we say that $P(x)$ holds $(s,p)$-quasi everywhere on $E$ (abbreviated as $q.e.$ on $E$).
\begin{defn}
A function $u : \Omega \rightarrow \R$ is said to be quasi-continuous (q.c.) if for every $\varepsilon > 0$, there exists a relatively open set $G \subset \Omega$ such that $\text{Cap}(s,p) (G) < \varepsilon$ and $u|_{\Omega \backslash G}$ is continuous.
\end{defn}
\begin{thm}[\cite{WARMA-15},Theorem 3.7]
	For every $u \in W^{s,p}_0(\Omega)$, there exists a unique quasi-continuous function $\tilde{u} : \Omega \rightarrow \mathbb{R}$ such that $\tilde{u} = u$ a.e. on $\Omega$.
\end{thm}

\begin{defn}
	$A\subset\Omega$ is considered an \emph{quasi open} set if for any $\varepsilon>0$ there exist $A^\varepsilon$ open set of $\Omega$ such that  Cap$(s,p)(A\triangle A^\varepsilon)\leq\varepsilon$, where $\triangle$ is the symmetric difference of sets.
\end{defn}
We represent the class of all $(s,p)$-quasi open subsets of $\Omega$ as $\A_{s,p}(\Omega)$, and for every $A\in \A_{s,p}(\Omega)$ we define the space
$$
W_0^{s, p}(A)=\{u\in W^{s, p}(\R^n)\, |\, u=0 \text{ q.e in }  \R^n\setminus A \}.
$$

$W_0^{s, p}(A)$ with norm inherited from $W^{s,p}_0(\Omega)$ is a closed subspace,therefore, is a reflexive Banach space. If $A$ is open the previous definition of $W^{s,p}_0(A)$ is equivalent to the usual one.

Following, we will provide the definition of a weak solution for \emph{$p$-Laplace} and
\emph{fractional $p$-Laplace} operator.
\begin{defn}
	Let \( \mathcal{B} \subset \mathbb{R}^n \) be a  bounded  set, given $s$ and $p$ such that \( 0 < s \leq 1 < p < \infty \), \(p'\) the Lebesgue conjugate of \(p\) and \( f \in L^{p'}(\mathcal{B}) \). We say that \( u \in W^{s,p}_0(\mathcal{B}) \) is a weak solution of 
	\begin{equation}\label{weak-solution}
	\begin{cases}
	(-\Delta_p)^s u = f \text{ in } \mathcal{B} \\
	u = 0 \text{ in } \mathbb{R}^n \setminus \mathcal{B}
	\end{cases}
	\end{equation} 
	if $u$ verifies 
	\[
	\langle (-\Delta_p)^s u , v\rangle = \int_\mathcal{B} f v \, dx,
	\]
	where 
	\begin{equation}
	\langle (-\Delta_p)^s u , v\rangle:=\begin{cases}\displaystyle\iint_{\mathcal{B} \times \mathcal{B}} \frac{|u(x) - u(y)|^{p-2}(u(x)-u(y))(v(x)-v(y))}{|x - y|^{n+sp}} \, dxdy,&\text{ if }s<1,\\
	\displaystyle\int_{\mathcal{B} } |\nabla u|^{p-2}uv \, dx&\text{ if }s=1,
	\end{cases}
	\end{equation}
	
	for every \( v \in W^{s,p}_0(\mathcal{B}) \).
\end{defn}
\begin{rem}
	If we consider the functional 
$ J_{s,p,\mathcal{B}} :W^{s,p}_0(\mathcal{B})\to  \R$ is defined as
	\begin{equation}
		J_{s,p,\mathcal{B}} (u)=
		\begin{cases} \displaystyle\iint_{\mathcal{B} \times \mathcal{B}} \frac{|u(x) - u(y)|^p}{|x - y|^{n+sp}} \, dx dy &\text{ if }s<1\\
			\displaystyle\int_{\mathcal{B}}|\nabla u|^p \,dx&\text{ if }s=1.
		\end{cases}	
	\end{equation}
	 $J_{s,p,\mathcal{B}}$ is a continuous functional with continuous Gateaux differential $J_{s,p,\mathcal{B}}'\colon W^{-s,p'}(\mathcal{B})\to W^{s,p}_0(\mathcal{B})$, given by
	$$
	\langle J_{s,p,\mathcal{B}}'(u),v\rangle=
	\begin{cases} \displaystyle\iint_{\mathcal{B} \times \mathcal{B}} \frac{|u(x) - u(y)|^{p-2}(u(x)-u(y))(v(x)-v(y))}{|x - y|^{n+sp}} \, dx dy &\text{ if }s<1,\\
		\displaystyle\int_{\mathcal{B}}|\nabla u|^{p-2}uv \,dx&\text{ if }s=1.
	\end{cases}	
	$$
\end{rem}	
	The monotonicity of the operators ensures, as in the standard case of an open set, the existence of weak solutions to \eqref{weak-solution}. Hence, given $A \in \A_{s,p}(\Omega)$, we denote by $u_A^s \in W_0^{s,p}(A)$ the unique weak solution to
\begin{equation}\label{def_u^a}
(-\Delta_p)^s v = 1 \text{ in } A, \quad v = 0 \text{ q.e. in } \mathbb{R}^n\setminus A.
\end{equation}
We will employ the following concept of set convergence, which is the fractional version of the $\gamma$-convergence of sets defined in \cite{DALMASO-93}:
\begin{defn}\label{def gamma}
	Let $\{A_k\}_{k\in\mathbb{N}} \subset \A_{s,p}(\Omega)$ and $A \in \A_{s,p}(\Omega)$. We say that $A_k \xrightarrow{\gamma} A$ if $u^s_{A_k} \to u_A^s$ in $L^p(\Omega)$.
\end{defn}
Now, we will show some properties related to $u_A^s$. 
\begin{lema}\label{lemma2}
	For any $A\in \A_{s,p}$ and $ u_A^s$ defined in \ref{def_u^a} the following holds:
	$$
	u_A^s=\max_{w\in W^{s,p}_0(\Omega)}\{ w\leq 0 \text{ in } \R^n\setminus A\text{ and } (-\Delta_p)^sw\leq 1 \text{ on } \Omega\}.
	$$
\end{lema}	
\begin{proof}	
	Let us define the convex set 
	$$	K_A=\{w\in W^{s,p}_0(\Omega)\colon w\leq 0 \text{ in } \R^n\setminus A	\},
	$$
	and let $w_A$ be the unique minimizer of $	I_s\colon K_A\to \R$ defined by
	$$
	I_s(w)=J_{s,p}(w)-\int_{\Omega} w\,dx.
	$$
	By direct calculus we have the following variational inequality
	\begin{equation}\label{variational_equation}
	\langle J_{s,p}'(w_A) , v-w_A\rangle 
	\geq \int_{\Omega} (v-w_A)\,dx\quad\forall v\in K_A.
	\end{equation}
	If we take $w_A^+$ as test function in \eqref{variational_equation}, we obtain that
	\begin{align*}
		0\leq\int_{\Omega}w_A^-\,dx \leq \langle J_{s,p}'(w_A) , w_A^-\rangle\leq
		-\iint_{(A\times A)^c}\frac{|w_A^-(x)-w_A^-(y)|^p}{|x-y|^{n+sp}}\,dx dy.
	\end{align*}
	From the preceding inequality, we deduce that $w_A^-=0$. Consequently, given that $w_A$ belongs to $K_A$, it follows that $w_A$ is in $W^{s,p}_0(A)$. Since $u_A^s$ represents the minimum of $I_s$ in $W^{s,p}_0(A)$, we can conclude that $w_A=u_A^s$.
	
	Next, given $v \in W^{s,p}_0(\Omega)$ such that $v \geq 0$, it follows that $-v \in K_A$. Hence, applying \eqref{variational_equation}
	\begin{align}\label{eq6}
		\langle J_{s,p}'( u_A^s) ,-v-  u_A^s\rangle=\langle J_{s,p}'( u_A^s) ,-v\rangle+\langle J_{s,p}'( u_A^s) ,-  u_A^s\rangle \geq\int_{\Omega} -v- u_A^s\,dx. 
	\end{align}
	Considering that 
	$$\langle J_{s,p}'( u_A^s) ,-  u_A^s\rangle = \displaystyle\int_{\Omega} - u_A^s\,dx,$$ the  inequality \eqref{eq6} can be expressed as
	$$
	\langle J_{s,p}'( u_A^s) ,v\rangle \leq\int_{\Omega} v\,dx.
	$$ 
	Since $v\in W^{s,p}_0(\Omega)$ is non negative but otherwise arbitrary, we get that $(-\Delta_p)^s u_A^s\leq 1$ in $\Omega$.
	
	Finally, if $w\leq 0$ in $\R^n \setminus A$ and $(-\Delta_p)^sw\leq 1$ in $\Omega$, then 
	$$
	(-\Delta_p)^sw\leq (-\Delta_p)^s u_A^s \text{ in }A\quad\text{ and } w\leq  u_A^s \text{ in }\R^n\setminus A.
	$$
	By comparison principle, $w\leq  u_A^s$ in $\R^n$, as we aim to demonstrate to conclude the lemma.
\end{proof}
To conclude this section, we will prove two lemmas that are crucial to our objective.
\begin{lema}\label{lemaA}
	Let $\{A_k\}_{k\in\N}$ be a sequence of \(quasi-open\) subsets of $\Omega$ such that $u_{A_k}^s$ converges weakly in $W^{s,p}_0(\Omega)$ to a function $u$, and let $\{v_k\}_{k\in \N}$ a sequence in $W^{s,p}_0(\Omega)$ such that $v_k=0$ q.e on $\Omega\setminus A_k$. Assume that $v_k\to v$ weakly in $W^{s,p}_0(\Omega)$. Then $v=0$ q.e. on $\{u=0\}$.
\end{lema}
Before presenting the proof, we will revisit the definition of $\Gamma$-convergence of functionals and some properties of the \emph{subdifferential set} of a convex function. See \cite{DALMASO-93} and \cite{BREZIS-LIBRO,MOTREANU14} respectively.
\begin{defn}
	Let $(X,d)$ be a metric space and $F_n,F\colon X\to \overline{\R}$, $n\in \N$. $F_n$ $\Gamma$-converge to $F$ if
	\begin{itemize}\label{liminfeq}
		\item For every $x\in X$ and for all sequence $\{x_n\}_{n\in\N}$ such that $x_n\to x$ hold that
		\begin{equation*}
			F(x)\leq\liminf_{n\to\infty}F_n(x_n).
		\end{equation*}
		\item For every $x\in X$, there exist a sequence $\{y_n\}_{n\in\N}$ such that $y_n\to y$ and
		\begin{equation*}\label{limsupeq}
			F(x)\geq\limsup_{n\to\infty}F_n(y_n).
		\end{equation*}
	\end{itemize}
\end{defn}
\begin{defn}
	Let $X$ a Banach space and $\phi\colon X\to\R \cup \{+\infty\}$ a convex function that is not identically $+\infty$. The \emph{effective domain} of the function $\phi$ is defined by
	$$
	Dom(\phi)=\{x\in X:\phi(x)<\infty\}.
	$$
	For $x\in Dom(\phi)$ the \emph{subdifferential} of  $\phi$ at $x$ is de set $\partial \phi(x)\subset X^*$ defined by
	\[
	\partial\phi(x)=\{x^*\in X^*\colon \phi(x+h)-\phi(x)\geq \langle x^*,h\rangle\text{ for all }h\in X\}
	\] 
\end{defn}
The following remark, we collect all properties that we will need.
\begin{rem}\label{rem1}Let us denote $\Gamma_0(X)$ the cone of convex, lower semicontinuous functions on a Banach space $X$ that are not identically $\infty$.
	\begin{itemize}
		\item	For every $x\in Dom(\partial\phi):=\{x\in X\colon \partial\phi(x)\not=\emptyset\}$ the set $\partial\phi(x)$  is nonempty, convex and $w^*-$closed. Moreover, if $\phi$ is convex and continuous at $x_0\in Dom(\phi)$, then $\partial \phi(x_0)$ is $w*-compact$.
		
		\item If $X$ is a reflexive Banach space, and $\phi\in\Gamma_0(X)$, then $Dom(\partial\phi)$ is dense in $Dom(\phi)$.
		
		\item	Finally, if $\phi^*(x^*)=\sup_{y\in X}\{\langle x^*,y \rangle-\phi(y)\}$ then
		$$x^*\in \partial \phi(x)\iff\phi(x)+\phi^*(x^*)=\langle x^*,x \rangle
		.$$ 
	\end{itemize}
\end{rem}

\begin{proof}[Proof of Lemma \ref{lemaA}]
	Let us define the functionals
	$$\Phi_k(w)=
	\begin{cases}
	J_{s,p,A_k}(w)  &\text{ if }w\in W^{s,p}_0(A_k)\\
	+\infty\qquad &\text{ otherwise}
	\end{cases}
	$$
	on the space $L^p(\Omega).$ From basic fact about the $\Gamma-$convergence there is a subsequence that $\Gamma$-converge to a functional $\Phi$ in $L^p(\Omega).$ The domain of $\Phi$, $Dom(\Phi)$, is contained in $L^p(\Omega)$ and as $\{v_k\}_{k\in\N}$ is a bounded sequence in $W^{s,p}_0(\Omega)$, we have
	$$
	\Phi(v)\leq\liminf_{k\to\infty}\Phi_k(v_k)\leq \liminf_{k\to\infty} J_{s,p}(v_k)<\infty,
	$$
	hence $v\in  Dom\,(\Phi)$. By Remark \ref{rem1}, the $Dom(\partial\Phi)$ is dense in $Dom(\Phi)$. Hence,  we will show that for every $\omega\in Dom(\partial\Phi)$, we have $\omega=0$ on $\{u=0\}$.
	
	Let us fix $\omega \in Dom(\partial\Phi)$ and $f\in\partial\Phi(\omega)$, by Remark \ref{rem1} $\omega$ is a minimum point of 
	$$
	\Psi(z)=\Phi(z)-\langle f,z\rangle.
	$$
	Let $\omega_k$ be the minimum of the functional
	$$
	\Psi_k(z)=\Phi_k(z)-\langle  f,z\rangle.
	$$
	Since $\Psi_k$ $\Gamma$-converges to $\Psi$ and Theorem 7.23 of \cite{DALMASO-93}, we have that $\{\omega_k\}_{k\in \N}\rightharpoonup \omega$ in $W^{s,p}_0(\Omega)$. Hence, we can define for every $\varepsilon>0$ the functional $ f_\varepsilon$ such that $\| f_\varepsilon-f\|_{L^p}<\varepsilon$, and $\omega_{k,\varepsilon}$ be the solution of the problem
	$$\begin{cases}
	(-\Delta_p)^s\omega_{k,\varepsilon}= f_{\varepsilon}&\quad\text{ in }A_k,\\
	\omega_{k,\varepsilon}=0 &\quad\text{ in } \Omega\setminus A_k.
	\end{cases}
	$$
	Then, from the well known inequality 
	\begin{align}
		(|a|^{p-2}a-|b|^{p-2}b)(a-b)\geq \begin{cases}
			c_1(|a|+|b|)^{p-2}|a-b|^2&\text{ if }1<p<2\\
			c_2|a-b|^p&\text{ if }2\leq p,
		\end{cases}
	\end{align}	
	where $c_1$ and $c_2$ are constants, we obtain $[\omega_{k,\varepsilon}-\omega_k]_{s,p}\leq c\| f_{\varepsilon}- f\|_{L^p(\Omega)}\leq c\varepsilon$ and (up to a subsequence) $\{\omega_{k,\varepsilon}\}$ converges weakly in $W^{s,p}_0(\Omega)$ to a function $\omega_\varepsilon$. Therefore, by lower semicontinuity of the norm, we get $[
	\omega_\varepsilon-\omega]_{s,p}\leq c\varepsilon$. Setting $c_\varepsilon=\| f_\varepsilon\|_{L^\infty(\Omega)}$, by the maximum principle we get $|\omega_{k,\varepsilon}|\leq c_\varepsilon u_{A_k}$ a.e in $\Omega$ for every $k\in\N$. Thus $|\omega_\varepsilon|\leq c_\varepsilon u$ a.e on $\Omega$. This implies that $\omega_\varepsilon=0$ on $\{u=0\}$.
\end{proof}

\begin{lema}\label{lemmaB}
	Let $A$ and $\{A_k\}_{k\in \N}$, be a quasi-open subsets of $\Omega$ such that $u^s_{A_k}\rightharpoonup u$ in $W^{s,p}_0(\Omega)$, with $u\leq  u_A^s$ on $\Omega$.
	Let $A^{\varepsilon}=\{ u_A^s>\varepsilon\}$ and assume that $u^s_{A_k\cup A^\varepsilon}$ converges to some function $u_\varepsilon$ weakly in $W^{s,p}_0(\Omega)$. Then, $u_\varepsilon\leq  u_A^s$ q.e. on $\Omega$.
\end{lema} 
\begin{proof}
	For every $\varepsilon>0$ let us define $v_\varepsilon=1-\frac{1}{\varepsilon}\min( u_A^s,\varepsilon)$, then we have $v_\varepsilon\in W_0^{s,p}(\Omega)$, $0\leq v_\varepsilon\leq 1$ on $\Omega$, $v_\varepsilon=0$ on $A^\varepsilon$ and $v_\varepsilon=1$ on $\Omega\setminus A.$ 
	
	Now set $w_k=\min(v_\varepsilon,u_{A_k\cup A^\varepsilon})$, then $w_k=0$ on $A^\varepsilon$ and $w_k=0$ on $\Omega\setminus(A_k\cup A^\varepsilon)$, so that $w_k=0$ on $\Omega\setminus A_k$. Moreover, $w_k$ converges to $\min(v_\varepsilon,u_\varepsilon)$. By Lemma \ref{lemaA} we get that $\min(v_\varepsilon,u_\varepsilon)=0$ on $\{u=0\}$. Therefore, $\min(v_\varepsilon,u_\varepsilon)=0$ on $\Omega\setminus A$. Since $v_\varepsilon=1$ on $\Omega\setminus A$, we have that $u_\varepsilon=0$ on $\Omega\setminus A$. 
	
	Finally, since $(-\Delta_p)^s u_{A_k\cup A^\varepsilon}\leq 1$ on $\Omega$, we have that $(-\Delta_p)^s u_\varepsilon\leq 1$ on $\Omega$, from Lemma \ref{lemma2}  $u_\varepsilon\leq  u_A^s$ on $\Omega$, and the proof is completed.
\end{proof}

\section{Statement and proof of the main result}\label{sec_teo1}
In this section, based in \cite{BUTTAZZO-DALMASO-93,BONDER-RITTORTO-18} and \cite{FUSCO-19} we will state and prove the main result.
\begin{thm}\label{teo1}
	Given $0 < s < 1<p<\infty $, $\Omega \subset \mathbb{R}^n$ be open and bounded set. Let $F_s : \A_{s,p}(\Omega) \to \R$ be such that
	\begin{itemize}
		\item[$(H_1)$] $F_s$ is lower semicontinuous with respect to the $\gamma$-convergence; that is, $A_k \xrightarrow{\gamma} A$ implies $F_s(A) \leq \liminf_{k\to\infty} F_s(A_k)$.
		\item[$(H_2)$] $F_s$ is decreasing with respect to set inclusion; that is, $F_s(A) \geq F_s(B)$ whenever $A \subset B$.
	\end{itemize}
	 Then, for every $0 < c < |\Omega|$, problem
	\begin{equation}\label{P}
		\min\{F_s(A): A \in \A_{s,p}(\Omega), |A| \leq c\},
	\end{equation} has a solution.
\end{thm}
Observe that there are numerous examples of functionals $F_s$ that satisfy the given hypotheses. For example, let $A\in \A_{s,p}(\Omega)$ be a admissible domain, and consider the eigenvalue problem:
$$
(-\Delta_\p)^su=\lambda^s u \text{ in }A\ \ ,u\in W^{s,p}_0(A).
$$
There, $\lambda^s\in\R$ is the eigenvalue parameter. It is well known, \cite{FRANZINA-14,BRASCO-PARINI-16,BRASCO-PARINI-16B} that there is a increasing discrete sequence  $\{\lambda_k^s(A)\}_{k\in\N}$ of eigenvalues. Therefore if we consider the application $A\to\lambda_k^s(A)$, the Theorem \ref{teo1} claims that for every $k \in \N$ and $0 < c < |\Omega|$, the minimum
\[
\min\{\lambda^s_{k}(A) : A \in \A_{s,p}(\Omega), |A| \leq c\}
\]
is achieved. 

Given the previous example, we can now proceed to the proof of Theorem \ref{teo1}.
\subsection{Proof of Theorem \ref{teo1}}
The sketch of argument is as follows, given $A\in \A_{s,p}$ we proof in Lemma \ref{lemma2} that $ u_A^s$ is the solution to
$$
\max\{w\in W^{s,p}_0(\Omega)\colon w\leq 0 \in \R^n\setminus A, (-\Delta_p)^sw\leq 1\}.
$$
Moreover $ u_A^s$ belong to 
\begin{equation}\label{K_S}
\K_s=\{w \in W_0^{s,p}(\Omega) : w \geq 0, (-\Delta_p)^s w \leq 1 \text{ in } \Omega\}
\end{equation}
 Then, we will define a functional $G$ on $\K_s$ satisfying
\begin{itemize}
	\item [$(G1)$]$G$ is decreasing on $\K_s$.
	\item [$(G2)$]$G$ is semicontinuous on $\K_s$ with the respect to the strong topology on $L^p(\Omega)$.
	\item [$(G3)$] $G( u_A^s)=F_s(A)$ for every $A\in \A_{s,p}$. 
\end{itemize}
to conclude that the problem 
$$
\min\{G(w):w\in \K_s,|\{w>0\}|<c\}
$$
has solution $w_0$. If we denote by $ A_0 = \{w_0 > 0\} $, then $ u \in A_0 $ is also a minimum point of $ G $ over the whole $ K_s $ subject to the condition $ |\{w > 0\}| \leq c $, and hence, $ A_0 $ is a minimizer for $ F_s $ in $ A_s(\Omega) $ subject to the condition $ |A| \leq c $.

First, we need to prove some properties about the set $\K_s$.
\begin{lema}\label{lemmaK}
	Given any sequence $\{u_k\}_{k\in\N}\subset \K_s$. Then, there exist $u\in K_s$ and a sub sequence $\{u_{k_j}\}_{j\in\N}$ such $u_{k_j}\to u$ in $L^p(\Omega)$. 
	
	Furthermore, for any $u, v$ belonging to $\K_s$, the maximum of $u$ and $v$, denoted as $\max\{u, v\}$, also belongs to $\K_s$.
\end{lema}
\begin{proof}	
	Given $ \{u_k\}_{k\in\N}\subset \K_s$, as each $u_k$ verifies $(-\Delta_p)^su_k\leq 1$ in $\Omega$ then for any non-negative test function $v\in W^{s,p}_0(\Omega)$ we have that 	
	$$
	\langle J'_{s,p}(u_k),v\rangle \leq \int_{\Omega}v\,dx.
	$$
	Taking $v:=u_k$, we deduce that $\{u_k\}_{k\in\N}$ is bounded in $W^{s,p}_0(\Omega)$, hence there exist a subsequence namely $\{u_k\}_{k\in\N}$, and $u\in W^{s,p}_0(\Omega)$ such that $u_k\rightharpoonup u$ in $ W^{s,p}_0(\Omega)$.
	
	Now, if we call
	 $$
	\eta_k(x,y)=\frac{|u_k(x)-u_k(y)|^{p-2}(u_k(x)-u_k(y))
}{|x-y|^{\frac{n+sp}{p'}}},	
	$$
	then $\eta_k$ is a bounded sequence in $L^{p'}(\R^n\times \R^n)$, where $p'$ is the Lebesgue conjugate of $p$. Therefore, there exist $\eta(x,y)$ $\in L^{p'}(\R^n \times \R^n)$ such that $\eta_k \rightharpoonup \eta$ weakly in $L^{p'}(\R^n \times \R^n)$. Therefore,
	\[
	\int_{\R^n \times \R^n} \eta_k(x, y) \frac{v(x) - v(y)}{|x - y|^{\frac{n}{p}+s}} \, dx \, dy \rightarrow \int_{\R^n \times \R^n} \eta(x, y) \frac{v(x) - v(y)}{|x - y|^{\frac{n}{p}+s}} \, dx \, dy.
	\]
	Since $u_k \rightarrow u$ almost everywhere in $\R^n$, one obtains that 
	$$\eta_k(x,y) \rightarrow \frac{|u(x) - u(y)|^{p-2}(u(x) - u(y))}{ |x-y|^{n+sp}}$$ a.e. in $\R^n \times \R^n$. Thus,
	\[
	\eta(x, y) = \frac{|u(x) - u(y)|^{p-2}(u(x) - u(y))}{|x - y|^{\frac{n+sp}{p'}}} .
	\]
Finally, for any $v\in W^{s,p}_0(\Omega)$ it follows that,
$$
		\langle J_{s,p}'(u),v\rangle =\lim_{k\to\infty}\langle J_{s,p}'(u_k),v\rangle\leq \int_{\Omega}v\,dx.
$$		
Hence, $u\in K_s$ and we conclude the fist part of the proof.

%
%

Let us denote $w=\max\{u,v\}$ and consider the convex set 
	$$
	E=\{z\in W^{s,p}_0(\Omega)\colon z\leq w \text{ in }\Omega\},
	$$
	and $z_0$ the unique minimizer of the functional $I\colon E\to \R$ defined by
	$$
	I(z)=J_{s,p}(z)-\int_{\Omega} z\,dx.
	$$
	Even more, $z_0$ verifies the variational inequality
	\begin{equation}\label{eq2}	
	\langle J_{s,p}'(z_0) , z-z_0\rangle 
	\geq \int_{\Omega} (z-z_0)\,dx\quad\forall v\in E.
	\end{equation}
	Let us see that $(-\Delta_p)^sz_0\leq 1.$ Given $\phi \in W^{s,p}_0(\Omega)$ such that $\phi\leq0$ and the funtional $i:\R_{+}\to\R$, defined as $i(t)=I(z_0+t\phi)$.	
	
	Observe that $i'(0)\geq 0$, in consequence, for any non-positive $\phi \in W^{s,p}_0(\Omega)$, it holds that 
	$$\langle J'_{s,p}(z_0), \phi\rangle \geq \int_{\Omega} \phi \,dx.$$
	 Then $$\langle J'_{s,p}(z), \phi\rangle \leq \int_{\Omega} \phi \,dx$$ for any $\phi \in W^{s,p}_0(\Omega)$, $\phi \geq 0$, and the claim follows.
	
	Now, we will prove that $z_0 \geq u$ (and for symmetry reasons that $z_0 \geq v$), from where it will follow that $z_0 \geq w$. Since $z_0 \in E$, the reverse inequality holds, and we can conclude that $z_0 = w \in \mathcal{K}_s$.
	Let $\eta=\max\{z_0,u\}$, since $\eta\in E$ we can use it as a test function in \eqref{eq2}. Thus,
	$$
	\langle J'_{s,p}(z_0),\eta-z_0\rangle\geq \int_{\R^n}(\eta-z_0)\, dx.
	$$
	Also, as $\eta-z_0\geq0$ and $(-\Delta_p)^su\leq 1$ in $\Omega$, it follows that
	$$
	\langle J'_{s,p}(u),\eta-z_0\rangle\leq \int_{\R^n}(\eta-z_0)\, dx.
	$$
	From both inequalities it is straightforward to see that
	$$
	0\leq\langle J'_{s,p}(z_0-u,\eta-z_0)\rangle\leq -2 J_{s,p}((u-z_0)^+)
	$$
	and then $(u - z_0)^+ = 0$ in $\R^n$, which implies that $z_0 \geq u$ in $\R^n$.	
\end{proof}
Now we can return to the proof of the theorem \ref{teo1}.
\begin{proof}
Let $F_s$ a functional that verifies hypothesis $(H_1)$ and $(H_2)$. Define the set 
$$
\A_{s}^c=\{A\in \A_{s}(\Omega)\colon |A|\leq c\}.
$$ 
Given $w\in \mathcal{K}_s$ define 
\begin{equation}
m(w)=\inf\{F(A)\colon A\in \A_{s}^c(\Omega), u_A^s\leq w\}.
\end{equation}
This functional is not lower semicontinuous, therefore we define $G$ to be the lower semicontinuous envelope of $m$ in $\mathcal{K}$ with the strong topology in $L^p$, i.e,
\begin{equation}\label{eq3}
G(w)=\inf\{\liminf_{k\to\infty} m(w_k)\},
\end{equation}
where the infimun is taken over all sequences $\{w_k\}_{k\in \N}$ in $\mathcal{K}$ such that $w_k\to w$ in $L^p(\Omega)$.

Let see us that $G$ is decreasing in $\K_s$. Let $u,v\in \K_s$ be such that $u\leq v$ and let $\{u_k\}_{k\in\N}\subset \K_s$ be such that $u_k\to u$ in $L^p(\Omega)$ and $m(u_k)\to G(u)$.
By Lemma \ref{lemmaK}, $v_k=\max\{v,u_k\}\in \K_s$ for each $k\in \N$ and $v_k\to v=\max\{v,u\}$ in $L^p(\Omega)$. Consequentely, since $m$ is non-increasing and $v_k\geq u_k$ for any $k\in\N$ we get 
$$
G(v)\leq \liminf_{k\to\infty}m(v_k)\leq \lim_{k\to\infty}m(u_k)=G(u).
$$
Now, let us see that $G$ verifies $(G_3)$, i.e.
\begin{equation}\label{G3}
G( u_A^s)=F(A),\text{ for every }A\in\A_{s,p}(\Omega).
\end{equation}
Given $A\in \A_{s,p}(\Omega)$, from \eqref{eq3} it follows that $G( u_A^s)\leq F(A)$. To prove the reverse inequality, it suffices to see that
$$
F(A)\leq \liminf_{k\to\infty}m(w_k)
$$
for any sequence $\{w_k\}_{k\in\N}\subset\K_s$ such that $w_k\to  u_A^s$ in $L^p(\Omega)$.
By definition of $m$, there exist $A_k\in \A_{s}^c(\Omega)$ such that 
$$
F(A_k)\leq m(w_k)+\frac{1}{k}\text{  and  }u^s_{A_k}\leq w_k.
$$ 
Observe that $u^s_{A_k}\in \K_s$ for each $k$, hence $\{u^s_{A_k}\}_{k\in\N}$ is bounded in $W^{s,p}_0(\Omega)$. Then, by \ref{lemmaK} up to a subsequence, there exist $u\in \K_s$ such that $u^s_{A_k}\to u$ in $L^p(\Omega)$.
Since $w_k\to  u_A^s$ in $L^p(\Omega)$, from $u^s_{A_k}\leq w_k$ we get $u\leq  u_A^s$.

Hence, given $\varepsilon>0$ we consider the set $A^\varepsilon=\{ u_A^s>\varepsilon\}$. Note that $u_{A_k\cup A^\varepsilon}^s\in \K_s$ and there exist $u^\varepsilon \in \K_s$ such that $ u_{A_k\cup A^\varepsilon}^s\to u^\varepsilon$ in $L^p(\Omega)$. Therefore, by Lemma \ref{lemmaB} $u^\varepsilon\leq  u_A^s$.

Next, let us see that $( u_A^s-\varepsilon)^+\leq  u_A^s.$ From
$$
( u_A^s-\varepsilon)^+(x)-( u_A^s-\varepsilon)^+(y)=\begin{cases}
 u_A^s(x)- u_A^s(y)&\quad\text{ if } x,y\in A^\varepsilon,\\
 u_A^s(x)-\varepsilon&\quad\text{ if } x\in A^\varepsilon\text{ and } y\not\in A^\varepsilon,\\
\varepsilon- u_A^s(y)&\quad\text{ if } y\in A^\varepsilon\text{ and } x\not\in A^\varepsilon,\\
0&otherwise,
\end{cases}
$$
for any $v\in W^{s,p}_0(A^\varepsilon)$ such that $v\geq 0$, we get that 
\begin{align*}
	\iint_{\R^n \times \R^n}&\frac{|(u(x)-\varepsilon)^+-(u(y)-\varepsilon)^+|^{p-2}((u(x)-\varepsilon)^+-(u(y)-\varepsilon)^+)(v(x)-v(y))}{|x-y|^{n+sp}}\geq\\&
	\iint_{\R^n \times \R^n}\frac{|u(x)-u(y)|^{p-2}(u(x)-u(y))(v(x)-v(y))}{|x-y|^{n+sp}}.
\end{align*}
Hence, $(-\Delta_p)^s( u_A^s-\varepsilon)^+\leq(-\Delta_p)^s  u_A^s=1=(-\Delta_p)^su^s_{A^\varepsilon}$ in $A^\varepsilon$. Moreover, since $0=( u_A^s-\varepsilon)^+=u^s_{A^\varepsilon}$ in $\R^n\setminus A^\varepsilon,$ from comparison principle, it follows that $( u_A^s-\varepsilon)^+\leq u_{A^\varepsilon}^s$ in $\R^n$.

In conclusion, we have obtained the following chain of inequalities
$$
( u_A^s-\varepsilon)^+\leq u_{A^\varepsilon}^s\leq u_{A_k\cup A^\varepsilon}^s.
$$
Since $u^\varepsilon\leq  u_A^s$ and $u_{A_k\cup A^\varepsilon}^s\to u^\varepsilon$ we conclude that
$$
( u_A^s-\varepsilon)^+\leq u^\varepsilon\leq  u_A^s.
$$
As the set $\{u^\varepsilon\}_{\varepsilon>0}$ is uniformly bounded in $W^{s,p}_0(\Omega)$, up to a subsequence we get that $u^\varepsilon\to  u_A^s$ in $L^p(\Omega)$. Hence by standard diagonal argument, there exist a sequence $\{\varepsilon_k\}_{k\in\N}$ such that 
$$
u_{A_k\cup A^{\varepsilon_k}} \to  u_A^s \text{ in }L^p(\Omega).
$$
By definition \ref{def gamma} $(A_k\cup A^{\varepsilon_k})$ $\gamma-$converges to $A$, obtaining the desired inequality,
$$
F(A)\leq \liminf_{k\to\infty}F(A_k\cup A^{\varepsilon_k})\leq \liminf_{k\to\infty}F(A_k)\leq \liminf_{k\to\infty}m(w_k).
$$
Now, let us focus on the final part of the proof.
Let $\{w_k\}\subset \K_s$ be a sequence such that $|\{w_k>0\}|\leq c$ and
$$
\lim_{k\to\infty}G(w_k)=\inf\{G(w)\colon w\in \K_s,\ |\{w_k>0\}|\leq c\}:=g.
$$
Also, there exist $w_0\in \K_s$ such that $w_k\to w_0$ in $L^{p}(\Omega)$ hence $|\{w_0>0\}|\leq c$, and since $G$ is semicontinuous on $\K_s$,
$$
g\leq G(w_0)\leq \liminf_{k\to\infty}G(w_k)=g
$$
so, $w_0$ verifies 
\begin{equation}\label{eq4}
w_0=\min\{G(w)\colon w\in \K_s,\ |\{w>0\}|\leq c\}.
\end{equation}
Finally, consider the set $A_0=\{w_0>0\}$. Therefore, $A_0\in \A_{s,p}^{c}(\Omega)$ and by Lemma \ref{lemma2}, $w_0\leq u^s_{A_0}$.
For every $A\in \A_{s,p}^c(\Omega)$, we know that $ u_A^s\in \K_s$, $|\{ u_A^s>0\}|\leq c$. Then, as $G$ is decreasing, verifies \eqref{G3} and \eqref{eq4} we have 
$$
F(A_0)=G(u^s_{A_0})\leq G(w_0)\leq G( u_A^s)=F(A).
$$
Therefore, $A_0$ is solution of the problem $\ref{P}$ and the proof is conclude.
\end{proof}

\section{Asymptotic behavior when $s\nearrow 1$}\label{sec_asint}
In this section, based on \cite{BBM-01,BONDER-SALORT-20} we will analyses the behavior when $s\nearrow 1$.

In order to perform such analysis we need to assume some asymptotic behavior
on the cost functional $F_s$. In order to do this, we need to define a notion of
convergence for sets when $s$ varies.
\begin{defn}
	Let $0<s_k\nearrow1$, and let $A_k\in\A_{s_k,p}(\Omega)$ and $A\in \A_{1,p}$. We say that, 
	$\A_k\xrightarrow{\gamma,k}\A$ if $u_{A_k}^{s_k}\to  u_A^1$ strongly in $L^p(\Omega)$.
	\end{defn}
\begin{thm}\label{teo2}
For any $0<s\leq1$, let $F_s\colon A_{s,p}(\Omega)\to \R$ be a functional that verifies
\begin{itemize}
	\item [($H_3$)] Continuity with respect to \(A\); that is, if \(A \in \A_1(\Omega)\), then
	$$
	F_1(A) = \lim_{s \uparrow 1} F_s(A). $$
	\item[($H_4$)]For every \(0 < s_k \uparrow 1\) and \(A_k \to A\), then
	\[ F_1(A) \leq \liminf_{k} F_{s_k}(A_k). \]
\end{itemize}

Assume moreover that $(H_1)$ and $(H_2)$ are satisfied. Then
$$
\min\{F_1(A):A\in \A_{1}(\Omega), |A|\leq c\}=\lim_{s\nearrow 1}\min\{F_s(A):A\in \A_{s}(\Omega), |A|\leq c\}.
$$
\end{thm}

%
%
%
Our primary aim is to establish that a sequence $\{u_k\}_{k\in\N}\in L^p(\Omega)$ such that $u_k\in \K_{s_k}$ is precompact and every accumulation point belong to $\K_1$.

\begin{lema}\label{lemma1} 
	Let $0<s_k\nearrow 1$ and let $u_k\in \K_{s_k}$. Then, there exist $u\in W^{1,p}_0(\Omega)$ and a subsequence $\{u_{k_j}\}_{j\in \N}$ such that $u_{k_j}\to u$ in $L^p(\Omega)$ and $u\in \K_1$, where 
	$$
	\K_1=\{w\in W^{1,p}_0(\Omega),\ w\geq0,\ (-\Delta_p)w\leq1\}.
	$$
	
\end{lema}
\begin{proof}
	As each $\K_{s_k}$ is a bounded set, and $W^{s_k,p}_0(\Omega)\subset W^{1,p}_0(\Omega)$ for all $k\in \N$. From \cite{BRASCO-PARINI-16} the optimal constant in Poincaré’s inequality
	$$
	 \|u\|_{L^p(\Omega)}^p \leq C(\Omega,p, s) [u]_{s,p}^p, 
	$$
	has a dependence on $s$ of the form	$ C(\Omega, p, s) \leq (1 - s)C(\Omega,p). $
	Therefore, there exists a constant $C$ such that
	$$
	\sup_{k\in \N}(1-s_k)[u_k]_{s_k,p}\leq C,
	$$
	and we can apply the well-known result [Theorem 4, \cite{BBM-01}] to conclude that there exist a subsequence $\{u_{k_j}\}_{k\in \N}$ such that $u_{k_j}\to u$ in $L^p(\Omega)$ and $u\in W^{1,p}_0(\Omega)$.
	 Finally, its clear that $u\geq0$ from results in  \cite{BONDER-SALORT-20,BRASCO-PARINI-16} and we get that $(-\Delta_p)u\leq1$ to conclude the proof.
	\end{proof}
Next, we want to analyze the continuity of \(  u_A^s\) as \(s\) approaches 1. 
\begin{lema}\label{lemma3}
	For every $A\in \A_{1}(\Omega)$, $u_A^s\to u_A$ strongly in $L^p(\Omega)$ when $s\nearrow 1$.	
\end{lema}
\begin{proof}
	Let us remind, that $u_A^s$ is the minimizer of  the functional 
	$$
I_s(w)=\begin{cases}
	J_{s,p,A}(w)-\int_{A} w\,dx&\textit{ if } w\in W^{s,p}_0(A),\\
	\infty&\text{ otherwise }.
\end{cases}.
	$$
	By \cite{PONCE-04}, we have that $J_{s,p,A}(w)$ $\Gamma$-converge to $J_{1,p,A}(w)$ in $L^p$. Hence, the minimizer of $J_{s,p,A}(w)-\int_{A} w\,dx$ converge to the minimizer of $J_{1,p,A}(w)-\int_{A} w\,dx$ strongly in $L^p(\Omega)$.
\end{proof}	
The following to lemmas, are the counterpart of Lemmas \ref{lemaA} and \ref{lemmaB}.
\begin{lema}\label{lemmaA2}
	Let $0<s_k\nearrow1$ and for every $k\in \N$ let $A_k\in \A_{s_k}(\Omega)$ be such that $u_{A_k}^{s_k}\to u$ strongly in $L^p(\Omega)$. Let $\{w_k\}_{k\in \N}\subset L^p(\Omega)$ be such that $w_k\in W^{s_k,p}_0(A_k)$ for every $k\in \N$ and $\sup_{k\in \N}(1-s_k)[w_k]_{s_k,p}^p<\infty$. Assume moreover that $w_k\to w$ strongly in $L^p(\Omega)$. Then, $w\in W^{1,p}_0(\{u>0\})$.
\end{lema}
\begin{proof}
	The proof closely resembles that of Lemma \ref{lemaA}, albeit with a few adjustments and the utilization of the compactness results from \cite{BBM-01}.
	\end{proof}
\begin{lema}\label{lemmaB2}
	Let $s_k\nearrow1$ and for every $k\in\N$ let $A_k\in \A_{s_k}(\Omega)$, $A\in\A_{1}(\Omega)$. Assume that $u_{A_k}^{s_k}\to u$ in $L^p(\Omega)$ and that $u\leq u_A$.
	Then, if $u_{A_k\cup A^\varepsilon}^{s_k}\to u^\varepsilon$ strongly in $L^p(\Omega)$, where $A^\varepsilon:=\{u_A>u\}$, its holds that $u^\varepsilon\leq u_A$.
\end{lema}
\begin{proof}
	By Lemma \ref{lemma1} we have that $u,u^\varepsilon\in W^{s,p}_0(\Omega)$. Let us define
	$$
	v^\varepsilon:=1-\frac{1}{\varepsilon}min\{u_A,\varepsilon\}.
	$$
	Observe that $0 \leq v^\varepsilon\leq 1$ and $v^\varepsilon=0$ in $A^\varepsilon$. 
	If we define
	\[u_{k,\varepsilon} := u_{,A_k^\varepsilon \cup A_\varepsilon}^s\qquad w_{k,\varepsilon} := \min\{v_\varepsilon, u_{k,\varepsilon}\},\]
	it holds that \(w_{k,\varepsilon} \geq 0\) since the comparison principle gives \(u_{k,\varepsilon} \geq 0\), and also \(v_\varepsilon \geq 0\).
	Since \(v_\varepsilon = 0\) in \(A_\varepsilon\), it follows that \(w_{k,\varepsilon} = 0\) in \(A_\varepsilon\). Moreover, since \(u_{k,\varepsilon} = 0\) in \(\mathbb{R}^n \setminus (A_k \cup A_\varepsilon)\), it holds that \(w_{k,\varepsilon} = 0\) in \(\mathbb{R}^n \setminus (A_k \cup A_\varepsilon)\), and consequently, \(w_{k,\varepsilon} \in W_0^{s_k,p}(A_k)\).
	Observe that \(w_{k,\varepsilon} \to w_\varepsilon := \min\{v_\varepsilon, u_\varepsilon\}\) strongly in \(L^p(\Omega)\), and then, applying Lemma \ref{lemmaB2}, we get \(w_\varepsilon \in W_0^{1,p}(\{u > 0\})\), from where \(w_\varepsilon = 0\) in \(\{u = 0\}\). The relation \(0 \leq u \leq u_A\) implies the inclusion \(\{u_A = 0\} \subset \{u = 0\}\), from where \(w_\varepsilon \in W_0^{1,p}(\{u_A > 0\})\). Moreover, since \(\{u_A > 0\} \subset A\), we have that \(w_\varepsilon = 0\) in \(\R^n \setminus A\). Now, being \(v_\varepsilon = 1\) in \(\mathbb{R}^n \setminus A\), we get \(u_\varepsilon = 0\) in \(\mathbb{R}^n \setminus A\), and in particular, \(u_\varepsilon \leq 0\) in \(\mathbb{R}^n \setminus A\).
	
	Let us show that \(-\Delta_p u_\varepsilon \leq 1\) in \(\Omega\). Observe that \(u_{k,\varepsilon} \in K_s^k\) and
	\(u_{k,\varepsilon} \to u_\varepsilon\) strongly in \(L^p(\Omega)\). Then, by Lemma \ref{lemma1}, \(u_\varepsilon \in K^1\) and \(-\Delta_p u_\varepsilon \leq 1\) in \(\Omega\). If we recall the Lemma \ref{lemma2} with $s=1$, 
	$$
	u_A=\max_{w\in W^{1,p}_0(\Omega)}\{ w\leq 0 \text{ in } \R^n\setminus A\text{ and } (-\Delta_p)w\leq 1 \text{ on } \Omega\},
	$$
	we conclude that $u^\varepsilon\leq u_A$.
\end{proof}	
\begin{prop}\label{prop1}
	Let $0 < s_k \uparrow 1$ , $A_k \in  \A_{s_k,p}^c(\Omega)$ be such that \(u^{s_k}_{A_k} \to u\) in \(L^p(\Omega)\). Then, there exist \(\tilde{A_k} \in \mathcal{A}_{s_k,p}(\Omega)\) such that \(A_k \subset \tilde{A_k} \) and \(\tilde{A_k}\) \(\gamma\)-converges to \(A := \{u > 0\}\).
	\end{prop}
\begin{proof}
Since $u^{s_k}_{A_k}\in \K_{s_k}$ and $u^{s_k}_{A_k}\to u$, by Lemma \ref{lemma1} $u\in \K_1$, and by Lemma \ref{lemma2}, $u\leq u_A$.

Consider $A^\varepsilon:=\{u_A> \varepsilon\}$ and observe that 
$$
u^{s_k}_{A^\varepsilon}\leq u^{s_k}_{A_k\cup A^\varepsilon}.
$$
Since $u^{s_k}_{A_k\cup A^\varepsilon}\in \K_{s_k}$, by Lemma \ref{lemma1}, there exist $u^\varepsilon \in W^{1,p}_0(\Omega)$ such that $u^{s_k}_{A_k\cup A^\varepsilon}\to u^\varepsilon$ in $L^p(\Omega)$.

Also, by Lemma \ref{lemma3}, \(u^{s_k}_{A^\varepsilon} \to u_{A^\varepsilon}\) strongly in \(L^p(\Omega)\). Then, we can pass to the limit as \(k \to \infty\) in the previous inequality to conclude that
\[u_{A^\varepsilon} \leq u_\varepsilon.\]
It can be easily checked that \(u_{A^\varepsilon} = (u_A - \varepsilon)^+\). From Lemma \ref{lemmaB2},
\[(u_A - \varepsilon)^+ \leq u_\varepsilon \leq u_A.\]
Thus, there exists a sequence \(0 < \varepsilon_k \downarrow 0\) such that
\[u^{s_k}_{{A_k} \cup A^{\varepsilon_k}} \to u_A\] strongly in \(L^p(\Omega)\).
That is, \(A_k \cup A_{\varepsilon_k} =: \tilde{A}_k\) \(\gamma\)-converges to \(A\).
\end{proof}
Now, we can give the proof of Theorem \ref{teo2}.
\begin{proof}
On one hand, by Theorem \ref{teo1} for each $k$ there exist $A_k\in \A_{s_k,p}^c(\Omega)$ such that 
$$
F_{s_k}(A_k)=\min\{F_{s_k}(A): A\in \A_{s_k,p}^c(\Omega)\}.
$$
Then if $A\in \A_{1,p}^c(\Omega)$ by condition $(H_1)$ we get that 
$$
\limsup_{k\to\infty}F_{s_k}(A_k)\leq \lim_{k\to\infty}F_{s_k}(A)=F_1(A).
$$
Hence,
\begin{equation}\label{eq_limsup}
\limsup_{k\to\infty} min\{F_{s_k}(A):A\in \A_{s_k,p}^c(\Omega)\}\leq  min\{F_{1}(A):A\in \A_{1,p}^c(\Omega)\}.
\end{equation}
On the other hand, by Lemma \ref{lemma1}, there is a $u\in W^{1,p}_0(\Omega)$ such that $u_{A_k}^{s_k}\to u$ in $L^p(\Omega)$. By Proposition \ref{prop1}, there exist $\tilde{A_k}\in \A_{s_k,p}^c(\Omega)$ such that $A_k\subset \tilde{A_k}$ and $\tilde{A_k}\gamma$-converge$ A:=\{u>0\}$. Since $u_{A_k}^{s_k}\to u$ then $|A|\leq c$.

From hypothesis $(H_2)$ and $(H_4)$ we conclude that 
$$
F_1(A)\leq \liminf_{k\to\infty}F_{s_k}(\tilde{A_k})\leq \liminf_{k\to\infty} F_{s_k}(A_k),	
$$
from where it follows that
\begin{equation}\label{eq_liminf}
\{F_{1}(A):A\in \A_{1,p}^c(\Omega)\}\leq \liminf_{k\to\infty} min\{F_{s_k}(A):A\in \A_{s_k,p}^c(\Omega)\}.
\end{equation}
Finally, combining \eqref{eq_limsup} and \eqref{eq_liminf} the proof is complete.
\end{proof}
\section{Anisotropic case}\label{sec_anis}
In this final section we will state the analogous concept of Theorem \ref{teo1} and its asymptotic behavior when we work in a anisotropic functional space.

Now, relying on \cite{CERESABONDER23}, let's begin by defining the anisotropic spaces that will be considered in this work.

Given $i=1,\dots, n$,  $s\in(0,1]$ and $p\in (1,\infty)$ we define the fractional $i^{th}-$Sobolev space as 
$$
W_i^{s,p}(\R^n)=\{u \in L^p(\R^n) \text{ such that } [u]_{s,p}^i< \infty\},
$$
where 
\[
[u]_{s,p,i} = 
\left \{
\begin{aligned}
&(1-s)s\int_{\R^n}\int_\R\frac{|u(x+he_i)-u(x)|^p}{|h|^{1+sp}}\, dh\, dx\ &\text{ if } \ &s < 1\\
&\frac{2}{p}\int_{\R^n} |\partial_{x_i} u|^p \, dx\ &\text{ if } \ &s=1.
\end{aligned}
\right .
\]
For convencience we are going to use the common notation 
\begin{align*}
	&\D_{s_i,p_i}(u):=\frac{|u(x+he_i)-u(x)|^{p_i}}{|h|^{1+s_ip_i}}, \\ &\D_{s_i,p_i}'(u):=\frac{|u(x+he_i)-u(x)|^{p_i-2}(u(x+he_i)-u(x))}{|h|^{1+s_ip_i}}.
\end{align*}
It is easy to see that $W^{s,p}_i(\R^n)$ is a Banach space with norm
$$
\|u\|_{s,p,i} = (\|u\|_p + [u]_{s,p,i})^{1/p},
$$
which is separable and reflexive. This space consists on $L^p$ functions that has some mild regularity in the $i^{th}-$variable. In order to control every variable, we proceed as follows. 
\begin{defn}\label{s,p}
	Given $\s=(s_1,\cdots,s_n)$ and $\p=(p_1,\cdots,p_n)$ we define the parameters
	$$
	\overline{\mathbf{s}}=\left(\frac{1}{n}\sum_{i=1}^n\frac{1}{s_i}\right)^{-1},\quad  \overline{\s\p}=\left(\frac{1}{n}\sum_{i=1}^n\frac{1}{s_ip_i}\right)^{-1}\text{ and }	
	\p^*_{\s} =\frac{n\overline{\s\p}/ \overline{\s}}{n-\overline{\s\p}}.
	$$
	Such that,
	\begin{enumerate}
		\item $0<s_i\leq 1$ for all $i\in \{1,\cdots n\}$.
		\item $1<p_1\leq \cdots\leq p_n<\infty$.
		\item $p_n<\p^*_\s$ and $\overline{\s\p}<n$.
	\end{enumerate}
	\end{defn}
Then, we define
$$
W^{\s,\p}(\R^n)=\bigcap_{i=1}^n W^{s_i,p_i}_i(\R^n).
$$
This anisotropic Sobolev space is then a separable and reflexive Banach space with norm
$$
\|u\|_{\s,\p} = \sum_{i=1}^n \|u\|_{s_i,p_i,i}.
$$
If we consider $\Omega \subset \R^n$, then we define the anisotropic Sobolev space of functions vanishing at $\R\setminus\Omega$ as
\begin{equation}\label{W0}
W^{\s,\p}_0(\Omega)=\{u \in W^{\s,\p}(\R^n) \colon u|_{\Omega^c}=0\}.
\end{equation}
By Poincaré inequality, the following proposition, we can use th norm 
$$
\|u\|_{\s,\p} := \sum_{i=1}^{n}( [u]_{s_i,p_i,i})^{1/{p_i}}.
$$
\begin{prop}
	Let $\Omega\subset\R^n$ be an bounded open subset. There exist a constant $C(\Omega,p)>0$ such that for all $u\in W^{s,p,i}_0(\Omega)$ holds that 
	$$
	\|u\|_{L^p(\Omega)}\leq C[u]_{s,p,i}.
	$$
	Hence, there exists a positive constant $C(\Omega, \p)>0$ such that for all $u \in W^{\s, \p}_0(\Omega)$, the following inequality holds for any $i \in 1, \ldots, n$.
	$$
	\|u\|_{L^{p_i}(\Omega)}\leq C\|u\|_{\s,\p}.
	$$
\end{prop}
\begin{proof}
	let $u$ be a function in $W^{s,p,i}_0(\Omega)$, we can assume that there exist $R>1$ such that $\supp{u}\subset Q_R = [-R,R]^n$. Hence,
	\begin{align*}
		[u]_{s,p,i}^{p}&=\int_{\R^n}\int_\R\frac{|u(x+he_i)-u(x)|^{p}}{|h|^{1+sp}}\,dh\,dx\\
		&\geq \int_{Q_R}\int_\R\frac{|u(x+he_i)-u(x)|^{p}}{|h|^{1+sp}}\,dh\,dx\\
		&\geq \int_{Q_R'}\int_{|x_i|\leq R}\int_{|x_i+he_i|\geq R}\frac{|u(x)|^{p}}{|h|^{1+sp}}\,dh\,dx_i\,dx'\\
		&\geq \int_{Q_R'}\int_{|x_i|\leq R}|u(x)|^{p}\int_{|h|\geq 2R}\frac{1}{|h|^{1+sp}}\,dh\,dx_i\,dx'\\
		&\geq \int_{Q_R'}\int_{|x_i|\leq R}|u(x)|^{p}\int_{|h|\geq 2R}\frac{1}{|h|^{1+p}}\,dh\,dx_i\,dx'\\
		&\geq C\|u\|_{p}^{p},
	\end{align*}
	where $Q_R' = [-R,R]^{n-1}$ and $dx' = dx_1\cdots,dx_{i-1},dx_{i+1}, dx_n$.
\end{proof}

Since we are dealing with admissible domains concerning the capacity of a set, we will define the analogous of $\text{Cap}(s,p)$, namely $\text{Cap}(\s,\p)$.
\begin{defn}
	Let $\Omega \subset \mathbb{R}^n$ be an open set, $\s$ and $\p$ defined in \ref{s,p}. Given $A \subset \Omega$. We define the Gagliardo $(\s,\p)$-capacity of $A$ relative to $\Omega$ as
	$$
	\text{Cap}(\s,\p)(A) = \inf \left\{ \|u\|_{\s,\p} : u \in C^\infty_c(\Omega), u \geq 0, A \subset \{u \geq 1\}^\circ \right\}.
	$$
\end{defn}
In this context, we say that a subset $A$ of $\Omega$ is a $(\s,\p)$-quasi open set if there exists a decreasing sequence $\{\omega_k\}_{k\in\mathbb{N}}$ of open subsets of $\Omega$ such that Cap$(\s,\p)(\omega_k) \to 0$, as $k \to +\infty$, and $A \cup \omega_k$ is an open set for all $k \in \mathbb{N}$.

We denote by $\A_{\s,\p}(\Omega)$ the class of all $(\s,\p)$-quasi open subsets of $\Omega$. And the notion of convergence will be defined as following.
\begin{defn} Let $\{A_k\}_{k\in\mathbb{N}} \subset \A_{\s,\p}(\Omega)$ and $A \in \A_{\s,\p}(\Omega)$. We say that $A_k \xrightarrow{\gamma_\s} A$ if $u_{A_k}^{\s} \to  u_A^{\s} \text{ in } L^{p_1}(\Omega).$ 
	
Moreover, if $\s_k:=(s_1^k,\cdots,s_n^k)\nearrow \vec{1}:=(1,\cdots,1)$. Given  $\{A_k\}_{k\in\mathbb{N}} \subset \A_{\s_k,\p}(\Omega)$ and  $\{A\}_{k\in\mathbb{N}} \in \A_{(1,\cdots,1),\p}(\Omega)$. We say that $A_k\xrightarrow{\gamma_1} A$ if $u_{A_k}^{s_k}\to u_A^{\vec{1}}$ in $L^{p_1}(\Omega)$.	
\end{defn}
\begin{defn}
	A function $u : \Omega \rightarrow \R$ is said to be $(\s,\p)-$quasi-continuous (q.c.) if for every $\varepsilon > 0$, there exists a relatively open set $G \subset \Omega$ such that $\text{Cap}(\s,\p) (G) < \varepsilon$ and $u|_{\Omega \backslash G}$ is continuous.
\end{defn}
The theorem presented below enables us to operate with the quasi-continuous representative of a function \(u\in W^{\s,\p}_0(\Omega)\). Its proof is similar to the proof of Theorem 3.7 in \cite{WARMA-15}, therefore we will omitted here.
\begin{thm}
	For every $u \in W^{\s,\p}_0(\Omega)$, there exists a unique relatively quasi-continuous function $\tilde{u} : \Omega \rightarrow \mathbb{R}$ such that $\tilde{u} = u$ a.e. on $\Omega$.
\end{thm}
Finally, we as wi will work with the fractional pseudo anisotropic p-laplacian defined in \eqref{fractiona pseudo p} we give the notion of weak solution.
\begin{defn}\label{weak-solution anis}
	Let \( \Omega \subset \R^n \) be a bounded  set, $\s$ and $\p$ defined in \ref{s,p}, $\p'$ is the coordinate Lebesgue conjugate of $\p$ and \( f \in L^{\p'}(\Omega) \). We say that \( u \in W^{\s,\p}_0(\Omega) \) is a weak solution of 
	\begin{equation}
	\begin{cases}
	(-\Delta_\p)^\s u = f \text{ in } \Omega \\
	u = 0 \text{ in } \R^n \setminus \Omega
	\end{cases}
	\end{equation} 
	if $u$ verifies 
	\[
	\langle (-\Delta_\p)^\s u , v\rangle = \int_\Omega f v \, dx,
	\]
	where 
	$$
	\langle (-\Delta_\p)^\s u , v\rangle:=\sum_{i=1}^{n}[(1-s_i)s_i]\iint_{\R^n \times \R} \D_{s_i,p_i}'(u)\, dhdx,
	$$
	for every \( v \in W^{\s,\p}_0(\Omega) \).
\end{defn}

As before, the monotonicity of the operators ensures, as in the standard case of an open set, the existence of weak solutions to \eqref{weak-solution anis}. Hence, given $A \in \A_{\s,\p}(\Omega)$, we denote by $u_A^{\s} \in W_0^{\s,\p}(A)$ the unique weak solution to
\begin{equation}
(-\Delta_\p)^\s v = 1 \text{ in } A, \quad v = 0 \text{ q.e. in } \mathbb{R}^n\setminus A.
\end{equation}

A potent tool associated with weak solutions is the comparison principle. While well-known in the isotropic case, in the anisotropic scenario, it necessitates stating and proving anew.
\begin{prop}[Comparison Principle]
	Let $u, v \in W^{\s,\p}_0(\Omega)$ be such that 
	$(-\Delta_{\p})^{\s}(u) \leq (-\Delta_{\p})^{\s}(v)$ in $\Omega$ in the sense that 
	\begin{align*}
		\sum_{i=1}^{n}\iint_{\R^n \times \R} \left[\D_{s_i,p_i}'(v)\right]&(\phi(x+he_i)-\phi(x))\,dh\,dx \\
		&\geq \sum_{i=1}^{n}\iint_{\R^n \times \R} \left[\D_{s_i,p_i}'(u)\right](\phi(x+he_i)-\phi(x))\,dh\,dx,
	\end{align*}
	in $\Omega$ whenever $\phi\in C^{\infty}_0(\Omega)$ and $\phi>0$. If $u \leq v$ in $\R^n\setminus \Omega$, then $u \leq v$ a.e. in $\Omega$.
\end{prop}

\begin{proof} The proof is based on \cite{LINDQVIST-14}, with a few changes.
	Let $u, v \in W^{\s, \p}_0(\Omega)$ be functions that verifies the hypothesis.
	Then for any $\phi\geq0$ in $W^{\s,\p}_0(\Omega)$ we have that
	\begin{equation}\label{eq1}
	\sum_{i=1}^{n}\iint_{\R^n \times \R}\left[\D_{s_i,p_i}'(v)-\D_{s_i,p_i}'(u)\right](\phi(x+he_i)-\phi(x))\,dh\,dx\geq0.
	\end{equation}
	We aim at showing that each integrand is non-positive for the choice $\phi(x)=(u-v)^+$. The identity
	$$
	|b|^{p-2}b-|a|^{p-2}a=(p-1)(b-a)\int_0^1|a+t(b-a)|^{p-2}\, dt.
	$$
	taking $a=u(x+he_i)-u(x)$ and $b=v(x+he_i)-v(x)$   then
	$$
	[\D_{s_i,p_i}'(v)-\D_{s_i,p_i}'(u)]=(p_i-1)\left[u(x)-v(x)-(u(x+he_i)-v(x+he_i))\right]Q_i(x,x+he_i)
	$$
	where 
	\begin{align*}
	Q_i(x,x+he_i)& =	\int_0^1|(u(x+he_i)-u(x))\\&+t\left((v(x+he_i)-v(x))-(u(x+he_i)-u(x))\right)|^{p_i-2}\, dt.
	\end{align*}
	Note that $Q_i\geq0$, and $Q_i=0$ if and only if $u(x+he_i)=u(x)$ and $v(x+he_i)=v(x)$. 
	We choose the test function $\phi=(u-v)^+$ and write $\psi=(u-v)$ each integrand of \eqref{eq1} is 
	$$(p_i-1)Q_i(x,x+he_i)|h|^{-1-s_ip_i}$$
	multiplied  with
	\begin{align*}
		&[\psi(x)-\psi(x+he_i)][\phi(x+he_i)-\phi(x)]=\\
		&-\left[\psi^+(x+he_i) - \psi^+(x)\right]^2 - \psi^-(x+he_i)\psi^+(x) - \psi^-(x)\psi^+(x+he_i).
	\end{align*}
	
	Hence each integrand contains only negative terms and, to avoid a contradiction, it is necessary that
	$$
	\psi^+(x+he_i)=\psi^+(x)\text{ or } Q_i(x,x+he_i)=0
	$$
	at a.e. point $(x,x+he_i)$. Therefore, the identity
	$$
	(u(x+he_i)-v(x+he_i))^+=(u(x)-v(x))^+
	$$
	must hold. Its follows that $u(x)-v(x)=C$ with $C$ constant, and by boundary conditions its follows that $C=0$ and $v\geq u$.
\end{proof}
To wrap up this endeavor, equipped with the properties and definitions of the anisotropic case, we are prepared to articulate the minimization problem. However, We will abstain from demonstrating Theorem \ref{teo1anis}, as it closely resembles the proofs of theorems in the preceding sections, albeit with slight modifications owing to our news lemmas and definitions.
\begin{thm}\label{teo1anis}
	Let $\s,\p$ defined in \ref{s,p}, and $\Omega \subset \R^n$ be open and bounded set. Let $F_\s : \A_{\s,\p}(\Omega) \to \R$ be such that 
	\begin{itemize}
		\item[$(H^{\s}_1)$] $F_\s$ is lower semicontinuous with respect to the $\gamma_\s$-convergence; that is, $A_k \xrightarrow{\gamma_\s} A$ implies $F_\s(A) \leq \liminf_{k\to\infty} F_\s(A_k)$.
		\item[$(H^{\s}_2)$] $F_\s$ is decreasing with respect to set inclusion; that is, $F_\s(A) \geq F_\s(B)$ whenever $A \subset B$.
	\end{itemize}
	are satisfied. Then, for every $0 < c < |\Omega|$, problem
	\begin{equation}\label{Panis}
	\min\{F_\s(A): A \in \A_{\s,\p}(\Omega), |A| \leq c\},
	\end{equation}
	has a solution. If $F_\s$ also verifies 
	\begin{itemize}
		\item [$(H^{\s}_3$)] Continuity with respect to \(A\); that is, if \(A \in \A_1(\Omega)\), then
		$$
		F_{\vec{1}}(A) = \lim_{\s \uparrow (1,\cdots,1)} F_\s(A). $$
		\item[$(H^{\s}_4)$]For every \( \s_k \uparrow (1,\cdots,1)\) and \(A_k \xrightarrow{\gamma_1} A\), then
		\[ F_1(A) \leq \liminf_{k\to \infty} F_{\s_k}(A_k). \]
	\end{itemize}
	Then
	$$
	\min\{F_1(A):A\in \A_{1}(\Omega), |A|\leq c\}=\lim_{\s\nearrow (1,\cdots,1)}\min\{F_\s(A):A\in \A_{\s,\p}(\Omega), |A|\leq c\}.
	$$
\end{thm}

As before, an example of a functional that satisfies the aforementioned hypotheses involves eigenvalues. In \cite{CERESABONDER24}, the existence of a sequence of eigenvalues related to the anisotropic eigenvalue of $(-\Delta_\p)^\s$ is proven in both the local and non-local cases.


\section*{Acknowledgements}
This work was supported by PIP No. 11220220100476CO. Ignacio Ceresa Dussel is a doctoral fellow of CONICET.
\bibliographystyle{plain}
\bibliography{References.bib}

\end{document}